OpenStudy (anonymous):

cosine of x divided by quantity one plus sine of x plus quantity one plus sine of x divided by cosine of x equals two times secant of x.

OpenStudy (freckles):

\[\frac{\cos(x)}{1+\sin(x)}+\frac{1+\sin(x)}{\cos(x)}=2\sec(x)?\]

OpenStudy (freckles):

assuming you are trying to show this is an identity I would first combine the fractions of the left

OpenStudy (freckles):

you know since we want to go from 2 terms to one term

OpenStudy (freckles):

need to make 2 terms one term by combining fractions

OpenStudy (anonymous):

rationalise the denominator of first term and proceed.

OpenStudy (anonymous):

\[\frac{ \cos x }{ 1+\sin x }\times \frac{ 1-\sin x }{ 1-\sin x }=\frac{ \cos x \left( 1-\sin x \right) }{ 1-\sin x }=?\]

OpenStudy (anonymous):

correction denominator \[=1-\sin ^2x\]